Database last updated on Tue, 03 Feb 2026 23:40:50 +0000 by localhost
Instrument constant changes for the week before 2023250 (3 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 19 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
dt
ed
fm
fz
hg
hp
rs
sl
uv
zs
250
RESOLUTE
#205
ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.161)
1
2
12
2
11
11
1
8
75
51
(0 sec)
Dead time (0 sec)
File name
Count
Serial
Expected (+/-5)
Measured
30-day history
B25023.205
1
#205
30
27